OpenStudy (anonymous):

what is the pattern: 1,4,9,16,25,.......

OpenStudy (psymon):

Usually a good idea is to subtract one number from the previous and see if it shows anything. 25-16 = 16 - 9 = 9 - 4 = 4 - 1 = Maybe youll see whats going on if youdo that.

OpenStudy (psymon):

Yep.

OpenStudy (anonymous):

first it goes by three and then u add two more each time to get a differnt number